OpenStudy (anonymous):

evaluate: integral from 3 to infinity: 1/(x-2)^(3/2) dx

OpenStudy (anonymous):

\[\int_3^\infty\frac{1}{(x-2)^{3/2}}\mathrm dx=\left[-\frac1{\frac32-1}\frac{1}{(x-2)^{3/2-1}}\right]_3^\infty\\ =0-(-2)=2 \]